OpenStudy (anonymous):

Write a recursive definition for the sequence 14, 10, 6, 2,… Common difference is 4 what do I do next?

OpenStudy (the_fizicx99):

f(n) = f(n – 1) + d \(\ \longrightarrow \) a_n = a_n-1 + d

OpenStudy (anonymous):

d is the common difference right? \[a_n = a_{n - 1} - 4\] How do we use this formula to find the next term tho

OpenStudy (the_fizicx99):

Well you know what a_1 is, it's 14 so to find a_2, \(\ \sf \Large a_2 = a_{2-1} - 4 \longrightarrow a_2 = a_1 - 4 \longrightarrow a_2 = 14 - 4 = a_2 = 10 \) Now look at your pattern, this is true, the second term or a_2 is 10!